# The thermodynamic structure and large-scale structure filament in MACS J0717.5+3745

### v2

@@ 5.1 Thermodynamical Maps
The temperature map on the top left of Fig. 4, shows an exceptionally hot cluster center, at around 24±$\pm 4$ keV, with a temperature dropping to 9.8±$\pm 0.8$ keV in the south and to 13±$\pm 3$ keV in the north-northwest. North of the temperature peak, and slightly west of the radio contours (see Fig. 9), we find a $\sim$ 20 keV elongated sub-structure; likewise, to the south and west of the peak, we see another $\sim$20 keV sub-structure.
Farther out towards the outskirts, in the north-northeast direction we see a spatially large region with temperatures $\sim$ 10 keV. To the west of MACS J0717.5+3745, after we pass the $\sim$ 10 keV region, we reach a seemingly different plasma distribution with a 5 keV plasma surrounded by 6-7 keV plasma. Results approaching the $R_{200}$ are probably biased by foreground structures and extended emissions to the west. In the south-southeast direction, we see a filamentary structure connecting the cluster with a group of galaxies. The filament temperature is reported in Tab. 5. In the map, the group itself has a temperature of 3.61±$\pm 0.17$ keV, whereas the temperature reported in Tab. 5 gives the group $R_{200}$ temperature. Based on previous lensing studies, it was determined that the filament continues farther out in this direction, even branching off into two separate filaments (Jauzac et al., 2018). Unfortunately, our X-ray data do not allow us to detect the X-ray emission associated with this part of the filamentary structure.
@@
The large pressure discontinuities that we observe in the center of the cluster imply relatively strong shocks. Assuming the Rankine-Hugoniot jump condition for temperature, the temperature jumps of $13\pm 2$ keV to $24\pm 4$ keV (northwards), and $10\pm 1$ keV to $21\pm 3$ keV (southwards), would correspond to Mach numbers of $\mathcal{M}=1.6\pm 0.4$ and $\mathcal{M}=1.9\pm 0.3$, respectively. van Weeren et al. (2017) gives an estimate of the Mach number of the shock derived from the radio spectral index of $\mathcal{M}=2.7$. Our estimate of the shock Mach number was derived from the thermodynamical maps directly, rather than from dedicated regions due to the complexity of the projected geometry. As this feature is composed of not one, but several bins, each larger than the PSF of XMM-Newton, using the map results should not contain extra uncertainty from spatial-spectral mixing. The difference between our respective measurements of the Mach number can be attributed to projection effects in the observation, the complex geometry of the system, as well as our flat, 1 Mpc line of sight assumption of for the density map. This line of sight depth should provide conservative estimates for the shock Mach number, as a larger line of sight depth would show a proportional decrease in the electron density, and would consequently result in a larger Mach number.
Both the pressure peak of the X-ray map and Sunyaev-Zeldovich peak are offset from the gravitational lensing peaks, but are also approximately in the middle of the two largest merging structures. The pressure and SZ signal peaks are also cospatial with the radio contours. Previous discussions about the nature of these radio structures proposed a hypothesis that these are radio relics, as indicated by the pressure jumps nearby, which are seen at some peculiar projected geometry (van Weeren et al., 2017).
@@
For a cluster in hydrostatic equilibrium, we expect to see an azimuthally symmetric and radially monotonously increasing entropy. Any departure from such a distribution is either due to the infall of subgroups hosting lower entropy material, uplift from the cluster center due to AGN feedback, or shock heating. The entropy map shows that the group in the southeast has a relatively low entropy near the center, which gradually increases with radius. The entropy of the X-ray bright bridge/filament appears lower than that of the surrounding ambient ICM at the same radius, which would be consistent with the potential infall of stripped material. At the position of the linear radio source, we see indications of an entropy discontinuity consistent with the presence of a shock front, where the lower entropy cluster center increases at the boundary. This linear bar region to the south of the cluster center is also cospatial with the temperature peak and associated X-ray shock features. Meanwhile, the characteristic ’V’ shape, which is present in the cluster center, seems to be associated with the colder region just north of the shock. The complex nature of the past subcluster merger activity and unknown geometry gives these shock discontinuities asymmetric and complex physical structures in projection.
@@
Curious features also become apparent in the various trend-divided thermodynamical maps. In the trend-divided density map, besides the filamentary region to the south-west, three subclusters near the cluster center show excess density, specifically the bar region associated with the shock front, the ’V’ shaped emission region at the cluster center, and the so-called ’finger’ region to the north-west of the cluster. These regions also appear as overpressure regions associated with entropy decrements in the trend-divided pressure and entropy maps. New interesting features, now visible in the trend-divided entropy map, are two high entropy regions originating from the cluster center and going to the northeast and southwest. Looking at the radio contours present in Fig. 9, the trend of this high entropy feature near the cluster center seems to trace the radio shock reported by van Weeren et al. (2017), however, near the outskirts of the cluster, it may be associated with the foreground structures present in the west of the cluster system. We note that the pressure excess associated with the shock is cospatial with the edge shown in the MUSTANG2 data and the kSZ and rSZ corrected MUSTANG2 data in Fig. 9; however, there is a bright radio galaxy in the foreground that makes any interpretation regarding this edge contain some degree of uncertainty.
@@ 5.2 The Filament
At the cluster redshift, the critical density of the Universe is $1.7\times 10^{-29}$ g cm-3. If we assume that the total baryon density is 0.044 of the critical density of the Universe and that the baryon mass fraction is 0.15 of the total mass density (Kirkman et al., 2003), we get an overdensity of our filament of 729 relative to the mean matter density of the Universe, and an overdensity of 401 relative to the critical density. Using the same volume, the baryonic mass of the filament is $\sim 6.1\times 10^{12}\leavevmode\nobreak\ \rm M_{\odot}$. Assuming that the baryon fraction of the filament equals the mean cosmic value, its total mass would be $\sim 7.6\times 10^{13}\leavevmode\nobreak\ \rm M_{\odot}$.
@@
Figure 10: Weak lensing map from the CLASH survey (Umetsu et al., 2015). The white box corresponds to the same FOV as the MUSTANG2 data and the other thermodynamical maps. The green box corresponds to the zoom-in region shown in Fig. 9.
Using the wide-field Subaru weak-lensing observations of Umetsu et al. (2014) to measure the mean surface mass density $\overline{\Sigma}$ within the filamentary region, we can provide an upper limit to the total mass of the filamentary structure. We note that the mass map shown in Fig. 10 was derived using Gaussian smoothing for visualization purposes (Umetsu et al., 2014), which is not suitable for accurate mass estimation. To this end, we employ the pixelized surface mass density map $\Sigma(\@vec{\theta}_{i})$ in the cluster field and its covariance matrix $C_{ij}$ from Umetsu et al. (2018), derived from the combination of two-dimensional reduced shear and azimuthally averaged magnification-bias constraints using the cluster lensing mass inversion 2D (clumi-2D) code. Combining complementary shear and magnification measurements, this method reconstructs the underlying surface mass density field around the lens, effectively breaking the mass-sheet degeneracy (see Umetsu, 2020, for a review of cluster weak lensing). The mass map $\Sigma(\@vec{\theta}_{i})$ covers a field of $24\times 24$ arcmin2 with $N_{\mathrm{pix}}=48\times 48$ pixels centered on the cluster. We find $\overline{\Sigma}=(3.7\pm 1.4)\times 10^{14}\leavevmode\nobreak\ \rm M_{\odot}$ Mpc-2 using an optimal (minimum variance) estimator, $\overline{\Sigma}=[A^{t}C^{-1}A]^{-1}A^{t}C^{-1}\Sigma$ (Appendix C of Umetsu et al., 2015), where $A$ is an $N_{\mathrm{pix}}\times 1$ mapping matrix whose elements $A_{i}$ are unity for those pixels lying inside the filamentary region and zero otherwise. The uncertainty for $\overline{\Sigma}$ is given by $\sigma(\overline{\Sigma})=(A^{t}C^{-1}A)^{-1/2}$. This translates into a projected mass inside the filament region of $(6.8\pm 2.7)\times 10^{13}\,\rm M_{\odot}$. These large uncertainties are reasonable, given that the weak lensing-based mass reconstruction has a pixel scale of 0.5 arcmin ($\sim 190$ kpc), which corresponds to only 7 pixels inside the filament region.
Taking these gas mass estimates ($4.2\times 10^{12}\leavevmode\nobreak\ \rm M_{\odot}$) and the projected total mass measurements ($6.8\pm 2.7\times 10^{13}\leavevmode\nobreak\ \rm M_{\odot}$) at face-value, the gas mass fraction of the filament is 4–10%. This range is comparable to the upper limit of 9% determined for the filament in Abell 222/223 (Dietrich et al., 2012) and within the range of 5–10% inferred for the filament system around Abell 2744 (Eckert et al., 2015b).
@@ 6 Summary
  • Thermodynamic maps produced from Chandra, XMM-Newton, and ROSAT data of MACS J0717.5+3745, using a new method for modelling both the astrophysical and instrumental backgrounds reveal a complicated ICM structure with several subsystems. The region south-southeast of the cluster center shows the presence of exceptionally hot 24±$\pm 4$ keV gas and pressure discontinuities in the cluster core indicate the presence of shocks. To the south-southeast, we detect the X-ray emission of a large-scale structure filament connecting the cluster with a group.
  • Trend-divided maps reveal that the localised density excess regions are related to the local overpressure regions, with a more complex entropy structure that appears approximately co-aligned to the spatial extent of the radio relic emission reported by van Weeren et al. (2017). The filament structure and other subclusters are more clearly visible after removing the cluster average trend for each physical quantity.
  • The temperature peak of 24±$\pm 4$ keV is also the pressure peak of the cluster and is spatially offset from the Sunyaev-Zeldovich peaks by 0.25 arcmin, around 100 kpc. We report shock Mach numbers of $\mathcal{M}=1.6\pm 0.4$ and $\mathcal{M}=1.9\pm 0.3$, for shocks located approximately halfway between the two lensing mass peaks.
  • Bayesian X-ray Analysis methods were used to disentangle different projected spectral signatures for the filament structure, with Akaike and Bayes criteria being used to select the most appropriate model to describe the various temperature components. The statistical measures indicate that the most complex model, involving CXB+Cluster+Group+Filament components is the best model to represent the filament structure. We report X-ray filament temperature of $3.1_{-0.3}^{+0.6}$ keV and density $3.8\pm 0.1\times 10^{-4}\,{\rm cm^{-3}}$, corresponding to an overdensity of the filament of 729 relative to the mean matter density of the Universe, and an overdensity of 401 relative to the critical density. We estimate the baryonic mass of the filament to be $\sim 6.1\times 10^{12}\leavevmode\nobreak\ \rm M_{\odot}$, while its total projected weak lensing measured mass is $\sim 6.8\pm 2.7\times 10^{13}\leavevmode\nobreak\ \rm M_{\odot}$, indicating a hot baryon fraction of 4–10%.
@@ Acknowledgements
Firstly, JPB would like to thank the reviewer for his/her comments to the first draft of the report. The implemented changes to the paper have added much to the scientific value and potential impact of the results. JPB would also like to give a big thank you to the members of the Chandra Ciao/Sherpa help desk and the various calibration and instrument scientists for all of their help and advice. Special thanks is to be given to Nick Lee, for spending a year and over 120 emails helping me troubleshoot the development of the instrumental background fitting procedure, and for helping understand the many discovered bugs regarding the AREASCAL keywords and the background scaling in Sherpa. Special thanks is also to be given to the XMM-Newton helpdesk team and the SAS developers, for their time and patience helping address the many SAS-related problems with the development of the analysis pipeline. I would like to thank Dominique Eckert, Jelle de Plaa, and Jelle Kaastra, for several discussions and method-related motivation; Peter Boorman and Johannes Buchner for their help with using BXA; and Michal Zajaček for many helpful discussions. The scientific results reported in this article are based in part on data obtained from the Chandra Data Archive. This research has made use of software provided by the Chandra X-ray Center (CXC) in the application packages Ciao and Sherpa; as well as NASA’s High Energy Astrophysics Software (HEASoft) packages Xspec. The scientific results reported in this article are also based in part on observations obtained with XMM-Newton, an ESA science mission with instruments and contributions directly funded by ESA Member States and NASA. JPB, NW, and TP acknowledge the financial support of the GAČR EXPRO grant No. 21-13491X. The material is based upon work supported by NASA under award number 80GSFC21M0002. K.U. acknowledges support from the National Science and Technology Council of Taiwan (grant NSTC 112-2112-M-001-027-MY3) and the Academia Sinica Investigator Award (grant AS-IA-112-M04). L.D.M. has been supported by the French government, through the UCAJ.E.D.I. Investments in the Future project managed by the National Research Agency (ANR) with the reference number ANR-15-IDEX-01.
@@ Appendix A Automatic Background Fitting
The caveat of this background fitting method, in particular, is that the background fitting procedure begins to fail when reaching very small regions with too few background counts. In other words, when the X-ray surface brightness is high (high S/N), the regions become small, which also increases the S/b and results in poor statistics. Coincidentally, this happens to be near the size of the PSF of each instrument. In these situations, the many parameters of the complex background model cannot be properly constrained. Therefore, the model becomes inappropriate with regard to the data quality of the background data. In these cases, we simplify the complex model to adjust to the data quality. We begin to do this by first removing the instrumental Gaussian lines so that we can fit a constant and the continuum. Then we try to also remove the continuum emission, and simply fit a constant model to the background. After each of these fits are performed, the reduced statistic, Akaike Information Criterion (AIC), and Bayes Information Criterion (BIC) are computed, and the most appropriate background model is selected for the dataset. For the very extreme cases, other background modelling methods and tools exist which could also be used; however, in this high S/b regime, a debate could be made about the need for a background model in the first place since the background can effectively be treated as noise (Buchner et al., 2014; Simmonds et al., 2018; Suzuki et al., 2021).
@@
[…]

### v3

@@ 5.1 Thermodynamical Maps
[…]
@@
The large temperature and pressure discontinuities that we observe in the center of the cluster indicate a potential presence of a shock candidate.
In particular, the pressure discontinuity to the southeast of the haloes C and D, (see Fig. 11) indicates that we may be looking at a shock candidate. This would also be consistent with the radio emission reported by van Weeren et al. (2017). Previous discussions about the nature of these radio structures in van Weeren et al. (2017) proposed the hypothesis that these are radio relics created as a result of subcluster merger activity, which are seen at some peculiar projected geometry. Based on our temperature map, we estimate the Mach number for the shock candidate somewhere between $\mathcal{M}=1.7\pm 0.3$ and $\mathcal{M}=2.0\pm 0.3$. van Weeren et al. (2017) gives an estimate of the Mach number of the shock derived from the radio spectral index of $\mathcal{M}=2.7$, which is not significantly different from our X-ray values, considering the various uncertainties. Alternatively, the observed pressure peak could be due to the gravitational potential peak of the two haloes (C and D). However, this scenario appears to be in tension with the observed steep temperature and pressure gradients.
@@
For a cluster in hydrostatic equilibrium, we expect to see an azimuthally symmetric and radially monotonously increasing entropy. Any departure from such a distribution is either due to the infall of subgroups hosting lower entropy material, uplift from the cluster center due to AGN feedback, or shock heating. The entropy map shows that the group in the southeast has a relatively low entropy near the center, which gradually increases with radius. The entropy of the X-ray bright bridge/filament appears lower than that of the surrounding ambient ICM at the same radius, which would be consistent with the potential infall of stripped material. At the position of the linear radio source, we see indications of an entropy discontinuity consistent with the presence of a shock front. This linear bar region to the south of the cluster center is also cospatial with the temperature/pressure peak and the associated X-ray shock features. Meanwhile, the characteristic ’V’ shape, which is present in the cluster center, seems to be associated with the colder region just north of the shock. The complex nature of the past subcluster merger activity and unknown geometry gives these shock discontinuities asymmetric and complex physical structures in projection.
@@
Curious features also become apparent in the various trend-divided thermodynamical maps. In the trend-divided density map, besides the filamentary region to the south-west, three subclusters near the cluster center show excess density, specifically the bar region associated with the shock front, the ’V’ shaped emission region at the cluster center, and the so-called ’finger’ region to the north-west of the cluster. These regions also appear overpressured and are associated with entropy decrements in the trend-divided pressure and entropy maps. New interesting features, now visible in the trend-divided entropy map, are two high entropy regions originating from the cluster center and going to the northeast and southwest. Looking at the radio contours present in Fig. 11, the trend of this high entropy feature near the cluster center seems to trace the radio shock reported by van Weeren et al. (2017), however, near the outskirts of the cluster, it may be associated with the foreground structures present in the west of the cluster system.
@@ 5.2 The Filament
At the cluster redshift, the critical density of the Universe is $1.7\times 10^{-29}$ g cm-3. If we assume that the total baryon density is 0.044 of the critical density of the Universe and that the baryon mass fraction is 0.15 of the total mass density (Kirkman et al., 2003), we get an overdensity of our filament of 729 relative to the mean matter density of the Universe, and an overdensity of 401 relative to the critical density. Using the same volume, the baryonic mass of the filament is $\sim 6.1\times 10^{12}\penalty 10000\ \rm M_{\odot}$. Assuming that the baryon fraction of the filament equals the mean cosmic value, its total mass would be $\sim 7.6\times 10^{13}\penalty 10000\ \rm M_{\odot}$.
@@
Figure 12: Weak lensing map from the CLASH survey (Umetsu et al., 2014). The white box corresponds to the same FOV as the MUSTANG2 data and the other thermodynamical maps. The green box corresponds to the zoom-in region shown in Fig. 11.
Using the wide-field Subaru weak-lensing observations of Umetsu et al. (2014) to measure the mean surface mass density $\overline{\Sigma}$ within the filamentary region, we can provide an upper limit to the total mass of the filamentary structure. We note that the mass map shown in Fig. 12 was derived using Gaussian smoothing for visualization purposes (Umetsu et al., 2014), which is not suitable for accurate mass estimation. To this end, we employ the pixelized surface mass density map $\Sigma(\@vec{\theta}_{i})$ in the cluster field and its covariance matrix $C_{ij}$ from Umetsu et al. (2018), derived from the combination of two-dimensional reduced shear and azimuthally averaged magnification-bias constraints using the cluster lensing mass inversion 2D (clumi-2D) code. Combining complementary shear and magnification measurements, this method reconstructs the underlying surface mass density field around the lens, effectively breaking the mass-sheet degeneracy (see Umetsu, 2020, for a review of cluster weak lensing). The mass map $\Sigma(\@vec{\theta}_{i})$ covers a field of $24\times 24$ arcmin2 with $N_{\mathrm{pix}}=48\times 48$ pixels centered on the cluster. We find $\overline{\Sigma}=(3.7\pm 1.4)\times 10^{14}\penalty 10000\ \rm M_{\odot}$ Mpc-2 using an optimal (minimum variance) estimator, $\overline{\Sigma}=[A^{t}C^{-1}A]^{-1}A^{t}C^{-1}\Sigma$ (Appendix C of Umetsu et al., 2015), where $A$ is an $N_{\mathrm{pix}}\times 1$ mapping matrix whose elements $A_{i}$ are unity for those pixels lying inside the filamentary region and zero otherwise. The uncertainty for $\overline{\Sigma}$ is given by $\sigma(\overline{\Sigma})=(A^{t}C^{-1}A)^{-1/2}$. This translates into a projected mass inside the filament region of $(6.8\pm 2.7)\times 10^{13}\,\rm M_{\odot}$. These large uncertainties are reasonable, given that the weak lensing-based mass reconstruction has a pixel scale of 0.5 arcmin ($\sim 190$ kpc), which corresponds to only 7 pixels inside the filament region.
Taking these gas mass estimates ($4.2\times 10^{12}\penalty 10000\ \rm M_{\odot}$) and the projected total mass measurements ($6.8\pm 2.7\times 10^{13}\penalty 10000\ \rm M_{\odot}$) at face-value, the gas mass fraction of the filament is 4–10%. This range is comparable to the upper limit of 9% determined for the filament in Abell 222/223 (Dietrich et al., 2012) and within the range of 5–10% inferred for the filament system around Abell 2744 (Eckert et al., 2015b).
@@ 6 Summary
[…]
  • We describe MACS J0717.5+3745 by creating a 3D cluster model assuming five spherical beta models. The geometric correction is based on the four strong lensing centers previously reported in the literature and the group at the end of the filament structure. Using this model, we calculate the effective line-of-sight length, which is used to calculate the volume to determine the density, pressure, and entropy in the thermodynamic maps.
  • Trend-divided maps reveal regions of excess density that are related to overpressured regions with complex entropy structure. These regions appear approximately co-aligned with the radio relic emission reported by van Weeren et al. (2017). The filament structure and other subclusters are more clearly visible after removing the cluster average trend for each physical quantity.
[…]
  • We present the first MUSTANG2 SZ maps of MACS J0717.5+3745, which were corrected for the kinetic and relativistic SZ effects using a model based on the best fit results from Adam et al. (2017b). The SZ maps show a close agreement with the X-ray pressure peak, and the kSZ model corrections are consistent with previously published work.
  • Bayesian X-ray Analysis methods were used to disentangle different projected spectral signatures for the filament structure, with Akaike and Bayes criteria being used to select the most appropriate model to describe the various temperature components. The statistical measures indicate that the most complex model, involving CXB+Cluster+Group+Filament components is the best model to represent the filament structure. We report X-ray filament temperature of $3.1_{-0.3}^{+0.6}$ keV and density $3.8\pm 0.1\times 10^{-4}\,{\rm cm^{-3}}$, corresponding to an overdensity of the filament of 729 relative to the mean matter density of the Universe, and an overdensity of 401 relative to the critical density. We estimate the baryonic mass of the filament to be $\sim 6.1\times 10^{12}\penalty 10000\ \rm M_{\odot}$, while its total projected weak lensing measured mass is $\sim 6.8\pm 2.7\times 10^{13}\penalty 10000\ \rm M_{\odot}$, indicating a hot baryon fraction of 4–10%.
@@ Acknowledgements
Firstly, JPB would like to thank the reviewer for his/her comments to the different drafts of the report. The implemented changes to the paper have added much to the scientific value and potential impact of the results. JPB would also like to give a big thank you to the members of the Chandra Ciao/Sherpa help desk and the various calibration and instrument scientists for all of their help and advice. Special thanks is to be given to Nick Lee, for spending a year and over 120 emails helping me troubleshoot the development of the instrumental background fitting procedure, and for helping understand the many discovered bugs regarding the AREASCAL keywords and the background scaling in Sherpa. Special thanks is also to be given to the XMM-Newton helpdesk team and the SAS developers, for their time and patience helping address the many SAS-related problems with the development of the analysis pipeline. I would like to thank Dominique Eckert, Jelle de Plaa, and Jelle Kaastra, for several discussions and method-related motivation; Peter Boorman and Johannes Buchner for their help with using BXA; and Michal Zajaček for many helpful discussions. The scientific results reported in this article are based in part on data obtained from the Chandra Data Archive. This research has made use of software provided by the Chandra X-ray Center (CXC) in the application packages Ciao and Sherpa; as well as NASA’s High Energy Astrophysics Software (HEASoft) packages Xspec. The scientific results reported in this article are also based in part on observations obtained with XMM-Newton, an ESA science mission with instruments and contributions directly funded by ESA Member States and NASA. JPB, NW, and TP acknowledge the financial support of the GAČR EXPRO grant No. 21-13491X. The material is based upon work supported by NASA under award number 80GSFC21M0002. K.U. acknowledges support from the National Science and Technology Council of Taiwan (grant NSTC 112-2112-M-001-027-MY3) and the Academia Sinica Investigator Award (grant AS-IA-112-M04). L.D.M. has been supported by the French government, through the UCAJ.E.D.I. Investments in the Future project managed by the National Research Agency (ANR) with the reference number ANR-15-IDEX-01.
@@ Appendix A Automatic Background Fitting
The caveat of this background fitting method, in particular, is that the background fitting procedure begins to fail when reaching very small regions with too few background counts. In other words, when the X-ray surface brightness is high (high S/N), the regions become small, which also increases the S/b and results in poor statistics. Coincidentally, this happens to be near the size of the PSF of each instrument. In these situations, the many parameters of the complex background model cannot be properly constrained. Therefore, the model becomes inappropriate with regard to the data quality of the background data. In these cases, we simplify the complex model to adjust to the data quality. We begin to do this by first removing the instrumental Gaussian lines so that we can fit a constant and the continuum. Then we try to also remove the continuum emission, and simply fit a constant model to the background. After each of these fits are performed, the reduced statistic, Akaike Information Criterion (AIC), and Bayes Information Criterion (BIC) are computed, and the most appropriate background model is selected for the dataset.
For the very extreme cases, other background modelling methods and tools exist which could also be used; however, in this high S/b regime, a debate could be made about the need for a background model in the first place since the background can effectively be treated as noise (Buchner et al., 2014; Simmonds et al., 2018; Suzuki et al., 2021).
@@
As evident by the two equations, the AIC and BIC both try to account for the goodness of fit between models, with a penalty given for the number of model parameters, where respectively, the AIC penalty is 2kk, versus the BIC penalty of $\ln(n)k$. Generally speaking, this punishes model complexity to prevent favouring models which overfit the data. These values can then be used to inform model selection; however, this is only applicable when comparing nested models relative to the same fitted data. The ‘cstat’ value is approximately equal to $1/2\ln(\mathcal{L})$ and has been shown to be a reliable estimator for model performance even with as few as 30 counts (Cash, 1979; Kaastra, 2017). We can then modify Eq. 11 to give metrics for our background model selection.